OpenStudy (anonymous):

Two garden plot are to have the same area. One is square and one is rectangular. The rectangular plot is 4 meters wide and 9 meters long. How large is one side of the square garden plot in meters?

OpenStudy (anonymous):

First find the area of rectangle as both the areas are equal, you can find the side of square then..

OpenStudy (anonymous):

\[Area \; of \; Rectangle = Length \times Width = 9 \times 4 = 36 \; m^2\]

OpenStudy (anonymous):

\[\text{Area Of Square} = \text{Area of Rectangle} \\ A_{\square} = 36 \; m^2\]

OpenStudy (anonymous):

So, \(Area \; of \; square = (side)^2\)

OpenStudy (anonymous):

I will say the rest

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

\[(side)^2 = 36 \\ \] Take square root both the sides, and find the \(side\) now..

OpenStudy (anonymous):

\[\huge \color{green}{\frak{Welcome \; \; To \; \; Openstudy... }}\]